an appliance cord is consist of two wires, about 2m long with individual wire which is about 3mm apart and carry a current of 8A DC. calculate the force exerted by the wires.

1
Expert's answer
2022-03-24T16:36:33-0400

Answer

Magnetic force

"F=\frac{2il^2}{r}\\=\frac{2*8*4}{0.003}\\=21.33*10^3N"